# LLM Agents for Education: Advances and Applications

### v1

@@ Appendix A Datasets & Benchmarks
Several datasets are designed to evaluate the pedagogical agents, such as ASSIST09 (Feng et al., 2009) and Junyi (Chang et al., 2015), which support knowledge tracing (KT) in K-12 math education, while others like EduAgent (Xu et al., 2024a) facilitate adaptive learning (AL) by dynamically adjusting content based on student profiles. In addition, error correction and detection (ECD) datasets, such as Virtual Teacher (Xu et al., 2024b) and MathCCS (Zhang et al., 2025b), assess LLM agents’ ability to identify and rectify student mistakes in math learning. Other datasets cater to writing, reading, and language learning, including FABRIC (Han et al., 2023), EssayJudge (Su et al., 2025), and EXCGEC (Ye et al., 2024b), which focus on feedback and generation (FCG) for student essays. MultiSim (Ryan et al., 2023) and Wang et al. (2024c) provide multi-lingual translation and storytelling benchmarks, expanding LLM capabilities beyond English-language education.
Several datasets support domain-specific educational agents across science, law, medicine, and computer science. Beyond their primary goal of evaluating pedagogical ability, these datasets assess LLM agents in domain-specific applications. They provide insights into how LLM agents can be adapted for specialized instruction, evaluating their ability to deliver subject-specific knowledge, facilitate problem-solving, and enhance interactive learning experiences across diverse educational fields. ScienceAgentBench (Chen et al., 2024a) and TheoremExplainBench (Ku et al., 2025) assess scientific reasoning and theorem explanation, while ML-Bench (Tang et al., 2023) and MLAgentBench (Huang et al., 2023a) focus on machine learning education. In law education, datasets like LawBench (Fei et al., 2023), LegalBench (Guha et al., 2023), and AgentCourt (Chen et al., 2024c) evaluate legal knowledge application, case analysis, and court simulations. Medical education datasets, including MedBench (Cai et al., 2024) and OmniMedVQA (Hu et al., 2024), test clinical reasoning and medical knowledge retrieval. For computer science, SWE-Bench (Yang et al., 2025) and Programming Feedback (Estévez-Ayres et al., 2024) assess code generation, debugging, and software engineering instruction. These benchmarks help refine LLM agents for specialized tutoring, enhancing AI-driven learning in professional fields.
Table 1. Summary of existing datasets and benchmarks of LLM agents for education.
@@
ASSIST09 KT Student Math K12 EN text 227k (Feng et al., 2009)
Junyi KT Student Math K12 ZH text 2.5M (Chang et al., 2015)
EduAgent AL & CS Student - Graduate EN text & image 1,015 (Xu et al., 2024a)
MathDial AL & KT Student Math K12 EN text 45 (Macina et al., 2023)
MultiArith AL Student Math K12 EN text 180 (Xu et al., 2024a)
CoMTA KT Student Math K12 EN text 153 (Scarlatos et al., 2025)
MaCKT KT Student Math K12 EN text 452 (Yang et al., 2024a)
Virtual Teacher ECD Student Math K12 ZH text & image 420 (Xu et al., 2024b)
MathCCS ECD Student Math K12 ZH text & image 420 (Zhang et al., 2025b)
MathTutorBench ECD & FCG Student Math K12 EN text 7 tasks (Macina et al., 2025)
MultiSim - Student Reading - Multi-lingual text 1.7M (Ryan et al., 2023)
FABRIC FCG Student Writing - EN text 1,782 (Han et al., 2023)
EssayJudge FCG Student Writing - EN text & image 1,054 (Su et al., 2025)
PROF FCG Teacher Writing - EN text 363 (Nair et al., 2024)
EXCGEC FCG Student Writing - ZH text 8,216 (Ye et al., 2024b)
Wang et al. (2024c) - All Translation - Multi-lingual text 70K (Wang et al., 2024c)
NewEpisode - Student Storytelling - EN text & image 24.5K (Wang et al., 2024e)
SD-Eval ECD Student Speaking - EN speech 7,303 (Ao et al., 2024)
@@
Programming Feedback FCG Teacher Computer Science - Code text 52 (Estévez-Ayres et al., 2024)
Review Critique FCG All Computer Science - EN text 440 (Du et al., 2024)
AAAR-1.0 FCG All Computer Science - EN text & image 1,000 (Lou et al., 2024)
ScienceAgentBench - All General Science - EN text 102 (Chen et al., 2024a)
TheoremExplainBench - All General Science - EN text 240 (Ku et al., 2025)
MLGym-Bench - All General Science - EN text 13 tasks (Nathani et al., 2025)
ML-Bench - All Machine Learning - EN text & image 9,641 (Tang et al., 2023)
MLAgentBench - All Machine Learning - EN text 13 tasks (Huang et al., 2023a)
SciCode - All General Science - EN text 338 (Tian et al., 2024a)
BLADE - All General Science - EN text 12 tasks (Gu et al., 2024)
DiscoveryBench - All General Science - EN text 1,167 (Majumder et al., 2024)
SUPER - All General Science - EN text 796 (Bogin et al., 2024)
E-EVAL - All Math & Language & General Science K12 ZH text 4,351 (Hou et al., 2024)
MedBench - All Medical - ZH text 40,041 (Cai et al., 2024)
CMB - All Medical - ZH text 280,839 (Wang et al., 2024a)
OmniMedVQA - All Medical - ZH text & image 127,995 (Hu et al., 2024)
MedEval - All Medical - EN text & image 22,779 (He et al., 2023)
OSWorld - All Computer Science - EN text & image 369 (Xie et al., 2024b)
Spider2-V - All Computer Science - EN text & image 812 (Cao et al., 2024)
VisualWebArena - All Computer Science - EN text & image 910 (Koh et al., 2024)
WebArena - All Computer Science - EN text & image 812 (Zhou et al., 2023)
SWE-Bench - All Computer Science - EN text 2,294 (Yang et al., 2025)
SWE-Bench M - All Computer Science - EN text & image 617 (Yang et al., 2024b)
Magentic-One - All Computer Science - EN text & image 617 (Fourney et al., 2024)
Lawbench - All Law - ZH text 20 tasks (Fei et al., 2023)
LegalBench - All Law - EN text 162 tasks (Guha et al., 2023)
LegalAgentBench - All Law - ZH text 300 tasks (Li et al., 2024a)
Agentcourt - All Law - ZH text 550 (Chen et al., 2024c)
SimuCourt - All Law - ZH text 420 (He et al., 2024a)

### v2

@@ Appendix B Datasets & Benchmarks
Several datasets are designed to evaluate the pedagogical agents, such as ASSIST09 Feng et al. (2009) and Junyi Chang et al. (2015), which support knowledge tracing (KT) in K-12 math education, while others like EduAgent Xu et al. (2024a) facilitate adaptive learning (AL) by dynamically adjusting content based on student profiles. In addition, error correction and detection (ECD) datasets, such as Virtual Teacher Xu et al. (2024b) and MathCCS Zhang et al. (2025d), assess LLM agents’ ability to identify and rectify student mistakes in math learning. Other datasets cater to writing, reading, and language learning, including FABRIC Han et al. (2023), EssayJudge Su et al. (2025), and EXCGEC Ye et al. (2024b), which focus on feedback and generation (FCG) for student essays. MultiSim Ryan et al. (2023) and Wang et al. (2024b) provide multi-lingual translation and storytelling benchmarks, expanding LLM capabilities beyond English-language education.
Several datasets support domain-specific educational agents across science, law, medicine, and computer science. Beyond their primary goal of evaluating pedagogical ability, these datasets assess LLM agents in domain-specific applications. They provide insights into how LLM agents can be adapted for specialized instruction, evaluating their ability to deliver subject-specific knowledge, facilitate problem-solving, and enhance interactive learning experiences across diverse educational fields. ScienceAgentBench Chen et al. (2024d) and TheoremExplainBench Ku et al. (2025) assess scientific reasoning and theorem explanation, while ML-Bench Tang et al. (2023) and MLAgentBench Huang et al. (2023c) focus on machine learning education. In law education, datasets like LawBench Fei et al. (2023), LegalBench Guha et al. (2023), and AgentCourt Chen et al. (2024b) evaluate legal knowledge application, case analysis, and court simulations. Medical education datasets, including MedBench Cai et al. (2024) and OmniMedVQA Hu et al. (2024), test clinical reasoning and medical knowledge retrieval. For computer science, SWE-Bench Yang et al. (2025) and Programming Feedback Estévez-Ayres et al. (2024) assess code generation, debugging, and software engineering instruction. These benchmarks help refine LLM agents for specialized tutoring, enhancing AI-driven learning in professional fields.
Table 2: Summary of existing datasets and benchmarks of LLM agents for education.
@@
ASSIST09 KT Student Math K12 EN text 227k Feng et al. (2009)
Junyi KT Student Math K12 ZH text 2.5M Chang et al. (2015)
EduAgent AL & CS Student - Graduate EN text & image 1,015 Xu et al. (2024a)
MathDial AL & KT Student Math K12 EN text 45 Macina et al. (2023)
MultiArith AL Student Math K12 EN text 180 Xu et al. (2024a)
CoMTA KT Student Math K12 EN text 153 Scarlatos et al. (2025)
MaCKT KT Student Math K12 EN text 452 Yang et al. (2024b)
Virtual Teacher ECD Student Math K12 ZH text & image 420 Xu et al. (2024b)
MathCCS ECD Student Math K12 ZH text & image 420 Zhang et al. (2025d)
MathTutorBench ECD & FCG Student Math K12 EN text 7 tasks Macina et al. (2025)
MultiSim - Student Reading - Multi-lingual text 1.7M Ryan et al. (2023)
FABRIC FCG Student Writing - EN text 1,782 Han et al. (2023)
EssayJudge FCG Student Writing - EN text & image 1,054 Su et al. (2025)
PROF FCG Teacher Writing - EN text 363 Nair et al. (2024)
EXCGEC FCG Student Writing - ZH text 8,216 Ye et al. (2024b)
Wang et al. (2024b) - All Translation - Multi-lingual text 70K Wang et al. (2024b)
NewEpisode - Student Storytelling - EN text & image 24.5K Wang et al. (2024e)
SD-Eval ECD Student Speaking - EN speech 7,303 Ao et al. (2024)
@@
Programming Feedback FCG Teacher Computer Science - Code text 52 Estévez-Ayres et al. (2024)
Review Critique FCG All Computer Science - EN text 440 Du et al. (2024)
AAAR-1.0 FCG All Computer Science - EN text & image 1,000 Lou et al. (2024)
ScienceAgentBench - All General Science - EN text 102 Chen et al. (2024d)
TheoremExplainBench - All General Science - EN text 240 Ku et al. (2025)
MLGym-Bench - All General Science - EN text 13 tasks Nathani et al. (2025)
ML-Bench - All Machine Learning - EN text & image 9,641 Tang et al. (2023)
MLAgentBench - All Machine Learning - EN text 13 tasks Huang et al. (2023c)
SciCode - All General Science - EN text 338 Tian et al. (2024a)
BLADE - All General Science - EN text 12 tasks Gu et al. (2024)
DiscoveryBench - All General Science - EN text 1,167 Majumder et al. (2024)
SUPER - All General Science - EN text 796 Bogin et al. (2024)
E-EVAL - All Math & Language & General Science K12 ZH text 4,351 Hou et al. (2024)
MedBench - All Medical - ZH text 40,041 Cai et al. (2024)
CMB - All Medical - ZH text 280,839 Wang et al. (2024d)
OmniMedVQA - All Medical - ZH text & image 127,995 Hu et al. (2024)
MedEval - All Medical - EN text & image 22,779 He et al. (2023)
OSWorld - All Computer Science - EN text & image 369 Xie et al. (2024b)
Spider2-V - All Computer Science - EN text & image 812 Cao et al. (2024)
VisualWebArena - All Computer Science - EN text & image 910 Koh et al. (2024)
WebArena - All Computer Science - EN text & image 812 Zhou et al. (2023)
SWE-Bench - All Computer Science - EN text 2,294 Yang et al. (2025)
SWE-Bench M - All Computer Science - EN text & image 617 Yang et al. (2024a)
Magentic-One - All Computer Science - EN text & image 617 Fourney et al. (2024)
Lawbench - All Law - ZH text 20 tasks Fei et al. (2023)
LegalBench - All Law - EN text 162 tasks Guha et al. (2023)
LegalAgentBench - All Law - ZH text 300 tasks Li et al. (2024a)
Agentcourt - All Law - ZH text 550 Chen et al. (2024b)
SimuCourt - All Law - ZH text 420 He et al. (2024b)